# With Ears to See and Eyes to Hear: Sound Symbolism Experiments with Multimodal Large Language Models

### v1

@@ 3.2 Results
4 Magnitude Symbolism
Figure 2: Examples of “Mil”-style (tiny) and “Mal”-style (huge) generations with DALL-E 3. In total, 50 images were generated, with 25 per condition (the entities remaining constant). The true ground truth is taken as the majority human vote.
4.1 Methodology
Image Dataset
Pseudowords
[…]
Task Setting
Models & Evaluation
We use the same models and evaluation protocols as in §3.1.
4.2 Results
Overall, similar to §3 we see mixed results in Table 2 as to which model performs best, but GPT-4 demonstrates higher levels of agreement across most conditions in a clearer pattern than what was seen in the Shape Symbolism experiment. In several cases, we see agreement that is significantly in line with human perception, with 90+agreement (e.g., GPT-4 in the Zil-Zal and Weel-Waal conditions). When comparing the Standard and informed prompts, we see a much more substantial increase in performance, indicating that the VLMs fundamentally understand the relationship between sound and perceived size, but were focussed on other aspects of the provided imagery when not explicitly directed towards size in the Standard condition. Regarding the LLaVA models, we see the mid-sized 13B variant outperforming the 7B and 34B models in most conditions (rather than performance increasing alongside parameter count).
@@ 4.2 Results
5 Iconicity Ratings
In this section we investigate whether LLMs demonstrate human-like associations between word forms and the entities/concepts they symbolise. [^6] We treat this as VLMs “imitating” an understanding of sound symbolism, as they of course cannot actually hear. Winter et al. (2023) present a dataset of 14k+ human judgements of iconicity to which we compare our LLM ratings (where, on a 7-point scale, “how” scored 1.3, whilst “woof” scored 6.8 due to being onomatopoeiac).
5.1 Methodology
Models
Existing work by Trott (2024) has investigated whether or not GPT-4 is able to reflect human judgements, and reported a Spearman correlation of 0.59. However, in trying to verify these findings, we see that our ratings differ quite strongly. For this reason, we re-run this experiment on the first 7k entries of Winter et al. (2023) and compare GPT-4 ratings as of late November 2023 with the human judgements in Winter et al. (2023) as well as the GPT-4 judgements from Trott (2024) that use an earlier version of GPT-4.777We test only on the first 7k entries due to cost. Whilst GPT-4-Turbo is markedly cheaper, this would not be a true reproduction. In the other cases, we use the entire 14,772 entry dataset, except for GPT-3.5-Turbo, where we remove 209 words that triggered OpenAI’s safeguarding filters.
Prompting
We adopt the same prompting strategy as Trott (2024). In doing so, we use a modified version of what was presented to human participants in Winter et al. (2023). In summary, we request ratings of iconicity on a 1-7 scale, where 1 is not at all iconic, and 7 is highly iconic. The full prompt presented to the models is available in Appendix C.2).
5.2 Results
Correlations between each LLM and human judgements are presented in Table 3. We observe that the ability of LLMs to rate the iconicity of English words appears to be dependent on model size. For instance, we see no true correlation with any of our FLAN T5 models (base = 250M, XL = 3B) or our smallest LLAMA-2 variant (7B), but we begin to see significant correlations with another 7B parameter model, that being instruction-tuned Mistral-7B. When investigating our larger models, we see the 13B LLAMA-2 variant demonstrate Spearman/Pearson correlations of .379 and .381, respectively. Interestingly, however, our largest LLAMA-2 model with 70B parameters performs worse at this task than the 13B variant as seen previously, with correlations of .304/.332. Regarding OpenAI models, GPT-3.5-Turbo demonstrates a moderate correlation with humans at .420/.439. Finally, GPT-4 presents the strongest correlations. However, we observe that (on the first 7k entries), the version of GPT-4 we use (late November 2023) performs worse than the earlier version presented by Trott (2024), further demonstrating how continuously updated models also require continuous evaluation.
Spearman Pearson
Model Corr. $p$-value Corr. $p$-value
FLAN-T5 Base -.035 <.001 -.037 <.001
FLAN-T5 XL .000 .991 -.002 .824
Mistral-7B-Instruct .382 <.001 .377 <.001
LLaMA-2 7B -.003 .687 .005 .544
LLaMA-2 13B .379 <.001 .381 <.001
LLaMA-2 70B .304 <.001 .332 <.001
GPT-3.5-Turbo .420 <.001 .439 <.001
GPT-4 (Trott, 2024) .575 <.001 .615 <.001
GPT-4 (Ours, Nov. ’23) .537 <.001 .594 <.001
Table 3: Correlations between human ratings from Winter et al. (2023) and LLMs. GPT-4 ratings were only collected for the first 7k examples due to cost. Trott (2024) report a Spearman correlation of .590 across the entire dataset.
6 Discussion
The Source of Sound Symbolism
Firstly, due to human languages exhibiting mostly regular orthography, auditory information in speech is moderately reflected in the spellings of words via grapheme sequences (a characteristic that has long been exploited by grapheme-to-phoneme conversion models). Through this, text-based models are able to learn associations between grapheme sequences and semantics, based on more abstract characteristics than morpheme combinations alone, such as phonaesthemes. Whilst such models have no embodied understanding of sound, such statistical patterns pose a viable signal for the implicit learning of sound-based phenomena.
[…]
As a result, the relatively weak performance of our tested models could also be explained by the relative lack of sound-symbolism-heavy language in the models’ training data which is overshadowed by more prosaic language forms that do not exploit these phenomena as readily. This in turn would explain why the closed-source models we tested (e.g. GPT-4/Gemini) outperform open-source models due to the significant (assumed) differences in parameter counts, allowing the larger closed models to retain more information regarding sound symbolism within the weights, in addition to being continuously updated with RLHF.
The results of our multimodal experiments additionally demonstrate that under certain conditions, VLMs show systematic disagreement with human labels, indicating the potential interference of additional knowledge contained within language model training data that influences the associations made between pseudowords and images that are not present in humans. However, it is important to note that in our experiments we compare language model selections against the majority vote or mean scores assigned by humans. Consequently, this results in a comparison to an ”ideal” human by necessity, overlooking individual differences in perception across humans (where for a decision to be ”human-like”, it has to match a choice made by any human, rather than the majority). Consequently, higher agreement levels can be observed when comparing to the choices of individual humans, as inter-human agreement is not perfectly aligned in these tests.
Future Directions
7 Conclusion
Limitations
Ethics Statement
Appendix A Appendix
A.1 Implementation Details
Kiki-Bouba & Mil-Mal
For the Kiki-Bouba and Mil-Mal sound symbolism experiments, we access gemini-pro-vision via the Google Gemini API. For GPT-4, we use gpt-4-vision-preview via the OpenAI Chat Completions API. For our open-source LLaVA models at various sizes, we specifically use llava-v1.6-vicuna-7b-hf, LLaVA-v1.6-vicuna-13b-hf and LLaVA-v1.6-vicuna-34b-hf from publicly available checkpoints on Hugging Face. We use default hyperparameters for all models to test their “off-the-shelf” capability.
Iconicity Ratings
Entities
@@ B.2 Results
The results of the soundscape description task can be seen in Table 4 and an example generation is presented in Figure 4 Overall, it can be seen that human evaluators thought positively of all 3 elements asked for from GPT-4, including the visual description (which would explain performance in the following section), soundscape description and onomatopoeia, with all criteria averaging at least 4. This therefore demonstrates that GPT-4 is able to provide convincing descriptions of auditory experiences when provided with a valid image. One key thing to note is that whilst the standard deviations are consistently low, onomatopoeia demonstrates the lowest consistently. This is to be expected when evaluating a literary device, as different people may have different preferences regarding onomatopoeia they would use for certain circumstances. Additionally, there may be cases where GPT-4 has described something such as a stream and assigned the onomatopoeia “whoosh”, which to one individual may sound too aggressive and resemble fast-moving water, when their own interpretation of a stream is more gentle (perhaps better suiting “lap lap”).

### v2

@@ A.2 Full Results
Shape Symbolism (Kiki/Bouba)
Table 2 presents the full unabridged results for the Sound Symbolism experiments presented in §3.
Magnitude Symbolism (Mil/Mal)
Table 3 presents the full unabridged results for the Magnitude Symbolism experiments presented in §4.
@@ B.2 Results
The results of the soundscape description task can be seen in Table 4 and an example generation is presented in Figure 7. Overall, it can be seen that human evaluators thought positively of all 3 elements asked for from GPT-4, including the visual description (which would explain performance in the following section), soundscape description and onomatopoeia, with all criteria averaging at least 4. This therefore demonstrates that GPT-4 is able to provide convincing descriptions of auditory experiences when provided with a valid image. One key thing to note is that whilst the standard deviations are consistently low, onomatopoeia demonstrates the lowest consistently. This is to be expected when evaluating a literary device, as different people may have different preferences regarding onomatopoeia they would use for certain circumstances. Additionally, there may be cases where GPT-4 has described something such as a stream and assigned the onomatopoeia “whoosh”, which to one individual may sound too aggressive and resemble fast-moving water, when their own interpretation of a stream is more gentle (perhaps better suiting “lap lap”).